OpenStudy (anonymous):

What is the value of x in the equation 8x-2y=48, when y=4 6 7 14 48

OpenStudy (asib1214):

7

Mehek (mehek14):

plug 4 in for y \(8x-2*4=48\\8x-8=48\\8x=56\\ \dfrac{56}{8}=x\)

Mehek (mehek14):

divide to get the answer

OpenStudy (asib1214):